The riverboat could go 48 miles downstream in the same time that it took her to go 32 miles upstream. If her speed was 10 mph in still water, what was the speed of the current?
Answer by scott8148(6628)

t = d / r ___ the times are the same (equal)

48 / (10 + c) = 32 / (10 - c)

"cross" multiplying ___ 480 - 48c = 320 + 32c